a rope is cut into three pieces in the ratio 1:3:5. given that the length of the longest side is 35m. a) calculate the length of the original rope
b) the length of the shortest side

Answer by stanbon(75887)

a rope is cut into three pieces in the ratio 1:3:5. given that the length of the longest side is 35m.
-----
Note:: 1:3:5 is the same as x:3x:5x
------
a) calculate the length of the original rope
5x = 35
x = 7
---
length = x + 3x + 5x = 9x = 9*7 = 63m
------------------------
b) the length of the shortest side
x = 7m
============
Cheers,
Stan H.